# Researcher Population Pyramids for Tracking Global Demographic and Gender Trajectories

### v1

@@ Acknowledgments
Country AUC F1 GAR AUC F1 GAR
Albania 0.932 0.913 0.813 0.997 0.998 0.892
Algeria 0.872 0.934 0.418 0.989 0.997 0.820
Australia 0.993 0.978 0.961 0.996 0.997 0.906
Austria 0.916 0.896 0.824 0.992 0.993 0.898
Bahrain 0.903 0.902 0.659 0.988 0.995 0.816
Belarus 0.862 0.911 0.431 0.905 0.941 0.685
Belgium 0.991 0.984 0.892 0.997 0.996 0.909
Bosnia and Herzegovina 0.960 0.952 0.759 0.986 0.990 0.898
Bulgaria 0.999 0.999 0.999 0.999 0.999 0.904
Canada 0.993 0.980 0.945 0.998 0.997 0.904
Cyprus 0.817 0.897 0.294 0.989 0.988 0.883
Czechia 0.953 0.957 0.710 0.968 0.973 0.814
Denmark 0.999 0.999 0.976 0.999 0.999 0.934
Egypt 0.891 0.900 0.625 0.988 0.995 0.823
Estonia 0.981 0.967 0.889 0.997 0.999 0.845
France 0.999 0.998 0.973 0.999 0.999 0.893
Germany 0.935 0.939 0.714 0.995 0.995 0.892
Ghana 0.904 0.921 0.647 0.992 0.991 0.879
Iceland 0.993 0.977 0.942 0.985 0.996 0.776
Iran, Islamic Republic of 0.860 0.898 0.512 0.996 0.998 0.867
Iraq 0.946 0.976 0.500 0.989 0.996 0.818
Ireland 0.999 0.993 0.977 0.999 0.999 0.943
Israel 0.939 0.933 0.740 0.987 0.991 0.827
Italy 0.995 0.990 0.987 0.999 0.998 0.926
Jamaica 0.913 0.927 0.609 0.991 0.990 0.875
Japan 0.931 0.910 0.792 0.946 0.971 0.676
Jordan 0.910 0.900 0.715 0.989 0.995 0.820
Kenya 0.882 0.900 0.581 0.990 0.989 0.882
Kuwait 0.890 0.908 0.608 0.989 0.996 0.822
Lebanon 0.895 0.897 0.691 0.990 0.997 0.817
Lesotho 0.845 0.908 0.428 0.991 0.990 0.880
Lithuania 0.955 0.931 0.921 0.982 0.989 0.834
Moldova, Republic of 0.879 0.913 0.549 0.907 0.939 0.674
Montenegro 0.885 0.928 0.476 0.994 0.993 0.899
Morocco 0.864 0.948 0.318 0.989 0.997 0.814
Netherlands 0.967 0.963 0.799 0.995 0.996 0.836
New Zealand 0.999 0.989 0.983 0.999 0.999 0.926
Nigeria 0.911 0.925 0.667 0.992 0.991 0.879
Norway 0.999 0.992 0.962 0.999 0.999 0.910
Oman 0.877 0.901 0.537 0.988 0.996 0.819
Papua New Guinea 0.942 0.922 0.830 0.991 0.990 0.884
Philippines 0.999 0.999 0.936 0.994 0.988 0.950
Poland 0.905 0.896 0.725 0.964 0.988 0.733
Portugal 0.972 0.959 0.858 0.999 0.999 0.962
Romania 0.938 0.946 0.592 0.988 0.988 0.864
Russian Federation 0.939 0.901 0.870 0.927 0.958 0.706
Saudi Arabia 0.892 0.903 0.598 0.990 0.998 0.788
Serbia 0.922 0.908 0.745 0.991 0.992 0.894
Somalia 0.943 0.902 0.857 0.988 0.995 0.821
South Africa 0.977 0.953 0.904 0.991 0.990 0.867
Spain 0.997 0.991 0.949 0.993 0.999 0.856
Sri Lanka 0.901 0.961 0.308 0.987 0.994 0.778
Sweden 0.998 0.994 0.951 0.998 0.999 0.920
Switzerland 0.998 0.990 0.963 0.999 0.999 0.929
Syrian Arab Republic 0.945 0.945 0.720 0.990 0.997 0.821
Tunisia 0.944 0.911 0.825 0.989 0.996 0.814
Turkey 0.981 0.968 0.891 0.994 0.993 0.839
United Arab Emirates 0.877 0.899 0.593 0.988 0.996 0.817
United Kingdom 0.995 0.984 0.938 0.998 0.998 0.880
United States 0.989 0.972 0.943 0.995 0.993 0.913
Zimbabwe 0.921 0.904 0.780 0.991 0.991 0.874
Table S7: Number of female and male authors in each country, along with the proportion of female authors. $n_{\text{female}}$: number of female authors. $n_{\text{male}}$: number of male authors. $p_{\text{female}}$: proportion of female authors.
Country $n_{\text{female}}$ $n_{\text{male}}$ $p_{\text{female}}$
Albania 3,802 2,644 59.0%
Algeria 25,220 30,153 45.5%
Australia 167,622 209,286 44.5%
Austria 46,871 70,258 40.0%
Bahrain 2,604 2,452 51.5%
Belarus 3,917 4,840 44.7%
Belgium 60,884 81,802 42.7%
Bosnia and Herzegovina 7,294 5,897 55.3%
Bulgaria 18,245 22,235 45.1%
Canada 248,601 288,415 46.3%
Cyprus 3,514 5,442 39.2%
Czechia 43,943 60,692 42.0%
Denmark 50,177 57,431 46.6%
Egypt 100,791 115,829 46.5%
Estonia 7,339 6,957 51.3%
France 293,981 451,893 39.4%
Germany 358,990 581,464 38.2%
Ghana 8,140 20,760 28.2%
Iceland 3,211 3,723 46.3%
Iran, Islamic Republic of 211,591 250,720 45.8%
Iraq 31,126 35,878 46.5%
Ireland 29,809 31,855 48.3%
Israel 53,115 72,756 42.2%
Italy 247,297 306,139 44.7%
Jamaica 1,370 1,179 53.7%
Japan 274,654 866,394 24.1%
Jordan 14,651 20,824 41.3%
Kenya 12,854 23,786 35.1%
Kuwait 4,597 7,363 38.4%
Lebanon 10,405 9,236 53.0%
Lesotho 236 292 44.7%
Lithuania 16,498 12,771 56.4%
Moldova, Republic of 3,779 3,446 52.3%
Montenegro 1,244 1,072 53.7%
Morocco 19,042 27,908 40.6%
Netherlands 114,796 171,909 40.0%
New Zealand 28,157 35,107 44.5%
Nigeria 49,250 109,430 31.0%
Norway 43,585 55,803 43.9%
Oman 4,891 5,150 48.7%
Papua New Guinea 570 1,145 33.2%
Philippines 31,406 27,892 53.0%
Poland 133,013 126,176 51.3%
Portugal 71,331 62,166 53.4%
Romania 50,551 40,920 55.3%
Russian Federation 188,225 281,228 40.1%
Saudi Arabia 53,278 72,306 42.4%
Serbia 25,834 23,279 52.6%
Somalia 249 670 27.1%
South Africa 47,444 58,705 44.7%
Spain 300,525 321,590 48.3%
Sri Lanka 4,607 6,640 41.0%
Sweden 73,425 89,455 45.1%
Switzerland 77,630 146,008 34.7%
Syrian Arab Republic 2,332 3,804 38.0%
Tunisia 23,287 19,589 54.3%
Turkey 187,101 212,756 46.8%
United Arab Emirates 10,680 12,530 46.0%
United Kingdom 423,058 537,556 44.0%
United States 2,185,469 2,812,284 43.7%
Zimbabwe 2,695 5,230 34.0%
Table S8: Estimated slope $\beta_{c}$ and coefficient of determination $R^{2}$ for each country.
Country $\beta_{c}$ $R^{2}$
Algeria 0.1537 0.9726
Egypt 0.1473 0.9938
Iraq 0.2615 0.9766
Jordan 0.1428 0.9933
Lebanon 0.1165 0.9839
Morocco 0.1063 0.9874
Saudi Arabia 0.1772 0.9797
Tunisia 0.1262 0.9254
United Arab Emirates 0.1556 0.9974
Australia 0.0572 0.9563
Canada 0.0502 0.9323
France 0.0418 0.9207
Germany 0.0519 0.9549
Italy 0.0576 0.9754
Japan 0.0192 0.9042
Spain 0.0696 0.9659
Sweden 0.0423 0.9521
United Kingdom 0.0473 0.9660
United States 0.0420 0.9679
Table S9: Estimated slope $\gamma_{c}$ and coefficient of determination $R^{2}$ for each country.
Country $\gamma_{c}$ $R^{2}$
Algeria 0.0075 0.9600
Egypt 0.0086 0.9721
Iraq 0.0075 0.8484
Jordan 0.0109 0.9489
Lebanon 0.0088 0.9515
Morocco 0.0029 0.6889
Saudi Arabia 0.0105 0.9145
Tunisia 0.0139 0.9760
United Arab Emirates 0.0101 0.9734
Australia 0.0069 0.9834
Canada 0.0071 0.9951
France 0.0061 0.9716
Germany 0.0074 0.9861
Italy 0.0048 0.9604
Japan 0.0041 0.9940
Spain 0.0065 0.9849
Sweden 0.0060 0.9800
United Kingdom 0.0066 0.9970
United States 0.0072 0.9956
Table S10: Estimated slopes before and after the Arab Spring ($\beta_{\text{before}}$ and $\beta_{\text{after}}$) for the log-transformed author counts in Egypt and Tunisia.
Country $\beta_{\text{before}}$ $\beta_{\text{after}}$ test statistic $p$-value
Egypt 0.055 0.054 0.447 0.661
Tunisia 0.092 0.021 22.069 $<0.0001$
Table S11: Estimated slopes before and after the Arab Spring ($\gamma_{\text{before}}$ and $\gamma_{\text{after}}$) for the proportion of female authors in Egypt and Tunisia.
Country $\gamma_{\text{before}}$ $\gamma_{\text{after}}$ test statistic $p$-value
Egypt 0.005 0.009 $-6.061$ $<0.0001$
Tunisia 0.014 0.009 3.187 0.006
Table S12: Inter-publication interval threshold (in years) by country and gender.
Country Female Male
Albania 9.04 10.59
Algeria 10.35 10.01
Australia 11.80 10.43
Austria 9.80 8.64
Bahrain 16.01 15.93
Belarus 12.84 13.01
Belgium 7.24 6.42
Bosnia and Herzegovina 10.33 10.45
Bulgaria 13.35 14.71
Canada 13.28 11.68
Cyprus 8.46 7.37
Czechia 8.68 8.59
Denmark 9.68 8.36
Egypt 14.68 14.93
Estonia 9.01 8.39
France 10.01 9.39
Germany 11.01 9.61
Ghana 10.57 10.01
Iceland 10.26 10.01
Iran, Islamic Republic of 7.72 7.01
Iraq 12.60 12.97
Ireland 15.69 13.59
Israel 11.10 9.96
Italy 8.76 7.92
Jamaica 25.86 24.29
Japan 12.01 10.47
Jordan 9.80 9.84
Kenya 11.26 11.49
Kuwait 13.26 13.24
Lebanon 13.07 12.89
Lithuania 7.01 7.42
Moldova, Republic of 11.01 11.52
Montenegro 11.50 11.02
Morocco 9.50 9.34
Netherlands 8.82 8.42
New Zealand 14.51 12.10
Nigeria 11.18 11.01
Norway 10.15 9.74
Oman 11.01 9.69
Philippines 15.36 15.01
Poland 9.01 8.91
Portugal 10.34 10.01
Romania 9.93 9.35
Russian Federation 11.24 12.01
Saudi Arabia 9.01 8.79
Serbia 10.01 10.15
South Africa 13.53 12.60
Spain 10.75 9.93
[…]

### v2

@@ Acknowledgments
Country AUC F1 $\text{GAR}_{\text{f}}$ $\text{GAR}_{\text{m}}$ AUC F1 $\text{GAR}_{\text{f}}$ $\text{GAR}_{\text{m}}$
Albania 0.934 0.918 0.781 0.842 0.997 0.998 0.916 0.855
Algeria 0.871 0.938 0.468 0.359 0.988 0.996 0.796 0.832
Australia 0.993 0.979 0.956 0.959 0.996 0.996 0.893 0.928
Austria 0.915 0.899 0.797 0.812 0.992 0.993 0.908 0.881
Bahrain 0.908 0.913 0.747 0.555 0.990 0.996 0.809 0.843
Belarus 0.859 0.907 0.439 0.414 0.902 0.936 0.708 0.648
Belgium 0.991 0.984 0.894 0.885 0.998 0.998 0.920 0.893
Bosnia and Herzegovina 0.962 0.954 0.762 0.781 0.986 0.991 0.904 0.882
Bulgaria 0.999 0.999 0.999 0.999 0.999 0.999 0.999 0.820
Canada 0.993 0.981 0.948 0.939 0.998 0.998 0.885 0.934
Cyprus 0.824 0.906 0.320 0.267 0.992 0.991 0.917 0.856
Czechia 0.959 0.965 0.768 0.670 0.970 0.975 0.847 0.783
Denmark 0.999 0.999 0.997 0.956 0.999 0.999 0.915 0.956
Egypt 0.895 0.904 0.649 0.601 0.988 0.996 0.814 0.834
Estonia 0.982 0.970 0.872 0.917 0.996 0.999 0.823 0.874
France 0.999 0.997 0.987 0.961 0.999 0.999 0.936 0.855
Germany 0.933 0.937 0.772 0.667 0.993 0.993 0.894 0.883
Ghana 0.906 0.923 0.673 0.590 0.992 0.989 0.914 0.854
Iceland 0.993 0.975 0.922 0.966 0.983 0.993 0.737 0.815
Iran, Islamic Republic of 0.857 0.892 0.526 0.516 0.996 0.998 0.856 0.893
Iraq 0.944 0.977 0.508 0.342 0.988 0.995 0.808 0.834
Ireland 0.999 0.993 0.983 0.976 0.999 0.999 0.930 0.958
Israel 0.937 0.933 0.734 0.759 0.988 0.991 0.852 0.800
Italy 0.995 0.991 0.985 0.988 0.999 0.999 0.946 0.910
Jamaica 0.909 0.929 0.608 0.576 0.992 0.990 0.907 0.852
Japan 0.928 0.908 0.759 0.811 0.945 0.970 0.625 0.726
Jordan 0.901 0.896 0.737 0.647 0.990 0.997 0.809 0.831
Kenya 0.886 0.906 0.516 0.633 0.992 0.991 0.910 0.847
Kuwait 0.888 0.901 0.655 0.556 0.990 0.996 0.808 0.836
Lebanon 0.897 0.906 0.693 0.597 0.990 0.996 0.807 0.838
Lesotho 0.846 0.908 0.375 0.474 0.992 0.994 0.089 0.259
Lithuania 0.953 0.927 0.909 0.942 0.982 0.991 0.837 0.829
Moldova, Republic of 0.871 0.908 0.559 0.524 0.896 0.933 0.689 0.638
Montenegro 0.888 0.932 0.439 0.526 0.993 0.994 0.912 0.888
Morocco 0.863 0.948 0.353 0.284 0.989 0.997 0.795 0.829
Netherlands 0.971 0.968 0.826 0.787 0.995 0.994 0.837 0.839
New Zealand 0.999 0.988 0.971 0.987 0.999 0.999 0.899 0.950
Nigeria 0.906 0.922 0.629 0.687 0.991 0.991 0.910 0.855
Norway 0.999 0.993 0.954 0.971 0.999 0.999 0.879 0.929
Oman 0.876 0.902 0.593 0.495 0.988 0.996 0.812 0.825
Papua New Guinea 0.939 0.918 0.839 0.824 0.992 0.990 0.911 0.854
Philippines 0.999 0.999 0.912 0.979 0.996 0.990 0.919 0.999
Poland 0.910 0.906 0.688 0.738 0.966 0.989 0.811 0.660
Portugal 0.974 0.961 0.864 0.862 0.999 0.999 0.956 0.961
Romania 0.931 0.939 0.591 0.566 0.987 0.987 0.906 0.830
Russian Federation 0.934 0.895 0.864 0.872 0.923 0.958 0.722 0.683
Saudi Arabia 0.888 0.903 0.645 0.488 0.987 0.996 0.768 0.799
Serbia 0.923 0.905 0.727 0.791 0.989 0.989 0.914 0.867
Somalia 0.945 0.975 0.551 0.400 0.989 0.996 0.803 0.837
South Africa 0.978 0.954 0.913 0.897 0.991 0.991 0.896 0.859
Spain 0.997 0.992 0.950 0.946 0.992 0.999 0.844 0.857
Sri Lanka 0.909 0.966 0.201 0.406 0.989 0.995 0.808 0.744
Sweden 0.998 0.993 0.946 0.953 0.999 0.999 0.920 0.918
Switzerland 0.997 0.988 0.955 0.964 0.999 0.999 0.943 0.910
Syrian Arab Republic 0.943 0.918 0.874 0.713 0.989 0.996 0.803 0.835
Tunisia 0.943 0.913 0.889 0.747 0.988 0.996 0.803 0.834
Türkiye 0.983 0.970 0.887 0.889 0.994 0.994 0.835 0.853
United Arab Emirates 0.878 0.906 0.603 0.544 0.990 0.997 0.816 0.832
United Kingdom 0.995 0.984 0.951 0.930 0.998 0.998 0.889 0.882
United States 0.989 0.972 0.952 0.932 0.995 0.993 0.887 0.928
Zimbabwe 0.923 0.906 0.789 0.780 0.992 0.990 0.912 0.852
Table S7: Number of authors by gender, proportion of female authors, and gender assignment rate for each country. NN: total number of authors with an assigned country. $n_{\text{female}}$: number of female authors. $n_{\text{male}}$: number of male authors. $p_{\text{female}}$: proportion of female authors among gender-assigned authors ($n_{\text{female}}/(n_{\text{female}}+n_{\text{male}})$). GAR: gender assignment rate $($$n_{\text{female}}+n_{\text{male}})/N$).
Country NN $n_{\text{female}}$ $n_{\text{male}}$ $p_{\text{female}}$ (%) GAR (%)
Albania 8,101 3,636 2,484 59.4% 75.5%
Algeria 82,032 24,186 28,860 45.6% 64.7%
Australia 449,116 157,027 193,939 44.7% 78.1%
Austria 140,857 43,527 64,650 40.2% 76.8%
Bahrain 6,383 2,377 2,163 52.4% 71.1%
Belarus 32,530 3,665 4,458 45.1% 25.0%
Belgium 193,359 56,615 75,125 43.0% 68.1%
Bosnia and Herzegovina 15,878 6,959 5,584 55.5% 79.0%
Bulgaria 46,690 17,182 20,643 45.4% 81.0%
Canada 709,352 232,752 266,578 46.6% 70.4%
Cyprus 13,567 3,162 4,876 39.3% 59.2%
Czechia 138,070 40,978 55,632 42.4% 70.0%
Denmark 137,867 47,737 54,154 46.9% 73.9%
Egypt 273,642 97,512 110,256 46.9% 75.9%
Estonia 18,136 6,826 6,224 52.3% 72.0%
France 975,619 280,487 429,656 39.5% 72.8%
Germany 1,362,026 342,417 551,829 38.3% 65.7%
Ghana 35,647 7,631 19,671 28.0% 76.6%
Iceland 7,445 2,977 3,382 46.8% 85.4%
Iran, Islamic Republic of 580,434 208,756 246,589 45.8% 78.4%
Iraq 115,842 25,416 27,660 47.9% 45.8%
Ireland 74,655 26,894 28,408 48.6% 74.1%
Israel 165,123 51,135 69,100 42.5% 72.8%
Italy 671,119 239,535 294,198 44.9% 79.5%
Jamaica 3,938 1,194 995 54.5% 55.6%
Japan 1,426,766 268,393 849,383 24.0% 78.3%
Jordan 41,137 13,926 19,738 41.4% 81.8%
Kenya 51,264 12,196 22,554 35.1% 67.8%
Kuwait 17,140 4,271 6,733 38.8% 64.2%
Lebanon 21,983 9,538 7,541 55.8% 77.7%
Lesotho 826 219 262 45.5% 58.2%
Lithuania 32,502 16,170 12,388 56.6% 87.9%
Moldova, Republic of 9,729 3,510 3,164 52.6% 68.6%
Montenegro 2,679 1,148 942 54.9% 78.0%
Morocco 77,995 17,994 26,155 40.8% 56.6%
Netherlands 394,621 107,770 160,882 40.1% 68.1%
New Zealand 75,098 25,651 31,850 44.6% 76.6%
Nigeria 261,895 47,770 106,339 31.0% 58.8%
Norway 121,078 41,343 52,325 44.1% 77.4%
Oman 15,633 4,635 4,749 49.4% 60.0%
Papua New Guinea 2,182 525 1,009 34.2% 70.3%
Philippines 79,309 30,165 26,658 53.1% 71.6%
Poland 359,952 129,861 121,841 51.6% 69.9%
Portugal 149,408 67,802 58,597 53.6% 84.6%
Romania 117,441 49,150 39,453 55.5% 75.4%
Russian Federation 940,103 180,330 257,350 41.2% 46.6%
Saudi Arabia 157,392 49,126 66,259 42.6% 73.3%
Serbia 57,701 25,065 22,165 53.1% 81.9%
Somalia 1,057 101 325 23.7% 40.3%
South Africa 141,129 45,024 55,139 45.0% 71.0%
Spain 801,513 291,516 310,020 48.5% 75.1%
Sri Lanka 43,124 4,210 6,172 40.6% 24.1%
Sweden 207,170 69,994 84,672 45.3% 74.7%
Switzerland 245,044 70,646 131,979 34.9% 82.7%
Syrian Arab Republic 8,132 2,525 3,694 40.6% 76.5%
Tunisia 62,780 22,157 18,328 54.7% 64.5%
Türkiye 440,360 183,318 208,001 46.8% 88.9%
United Arab Emirates 34,293 9,507 10,544 47.4% 58.5%
United Kingdom 1,282,607 389,777 490,191 44.3% 68.6%
United States 6,560,519 2,119,214 2,711,315 43.9% 73.6%
Zimbabwe 10,497 2,495 4,781 34.3% 69.3%
Table S8: Estimated slope ($\beta_{c}$) and coefficient of determination ($R^{2}$) for each country. Standard errors (SEs) and $p$-values are computed using a Newey-West estimator to account for potential autocorrelation.
Country $\beta_{c}$ $R^{2}$ SE $p$-value
Algeria 0.1537 0.9726 0.0082 < 0.0001
Egypt 0.1473 0.9938 0.0031 < 0.0001
Iraq 0.2613 0.9769 0.0099 < 0.0001
Jordan 0.1429 0.9933 0.0021 < 0.0001
Lebanon 0.1210 0.9829 0.0053 < 0.0001
Morocco 0.1063 0.9874 0.0032 < 0.0001
Saudi Arabia 0.1771 0.9800 0.0067 < 0.0001
Tunisia 0.1262 0.9253 0.0115 < 0.0001
United Arab Emirates 0.1569 0.9973 0.0015 < 0.0001
Australia 0.0572 0.9564 0.0040 < 0.0001
Canada 0.0502 0.9324 0.0045 < 0.0001
France 0.0418 0.9207 0.0039 < 0.0001
Germany 0.0519 0.9549 0.0036 < 0.0001
Italy 0.0576 0.9754 0.0030 < 0.0001
Japan 0.0192 0.9042 0.0022 < 0.0001
Spain 0.0696 0.9659 0.0041 < 0.0001
Sweden 0.0423 0.9521 0.0032 < 0.0001
United Kingdom 0.0473 0.9660 0.0029 < 0.0001
United States 0.0420 0.9678 0.0026 < 0.0001
Table S9: Estimated slope ($\gamma_{c}$) and coefficient of determination ($R^{2}$) for each country. Standard errors (SEs) and $p$-values are computed using a Newey-West estimator to account for potential autocorrelation.
Country $\gamma_{c}$ $R^{2}$ SE $p$-value
Algeria 0.0075 0.9598 0.0004 < 0.0001
Egypt 0.0086 0.9719 0.0003 < 0.0001
Iraq 0.0082 0.8495 0.0007 < 0.0001
Jordan 0.0110 0.9487 0.0006 < 0.0001
Lebanon 0.0075 0.9048 0.0007 < 0.0001
Morocco 0.0029 0.6891 0.0005 < 0.0001
Saudi Arabia 0.0106 0.9138 0.0010 < 0.0001
Tunisia 0.0139 0.9759 0.0006 < 0.0001
United Arab Emirates 0.0098 0.9700 0.0003 < 0.0001
Australia 0.0069 0.9833 0.0003 < 0.0001
Canada 0.0071 0.9951 0.0001 < 0.0001
France 0.0061 0.9716 0.0003 < 0.0001
Germany 0.0074 0.9861 0.0003 < 0.0001
Italy 0.0048 0.9604 0.0003 < 0.0001
Japan 0.0041 0.9940 0.0001 < 0.0001
Spain 0.0065 0.9849 0.0002 < 0.0001
Sweden 0.0060 0.9800 0.0002 < 0.0001
United Kingdom 0.0066 0.9970 0.0001 < 0.0001
United States 0.0072 0.9956 0.0001 < 0.0001
Table S10: Inter-publication interval thresholds (in years) for female and male authors when the survival probability threshold is set to 1%, 2%, and 5%.
1% 2% 5%
Country Female Male Female Male Female Male
Albania 8.18 9.88 6.92 7.09 4.67 4.33
Algeria 10.12 9.50 7.75 6.68 5.01 3.98
Australia 11.93 9.26 6.88 5.26 3.08 2.36
Austria 10.40 7.85 6.16 4.52 3.02 2.09
Bahrain 13.93 13.80 9.76 8.53 5.09 4.01
Belarus 11.52 11.27 8.46 7.45 4.70 3.83
Belgium 7.59 5.83 4.40 3.25 2.09 1.58
Bosnia and Herzegovina 10.01 10.01 7.01 6.93 4.05 4.01
Bulgaria 12.01 13.64 7.93 8.56 4.01 4.07
Canada 13.76 10.51 8.20 6.17 3.97 2.92
Cyprus 8.76 6.26 5.79 4.09 3.22 2.16
Czechia 8.01 7.92 5.16 4.68 2.84 2.25
Denmark 10.61 7.92 6.10 4.50 3.00 2.01
Egypt 13.75 14.01 8.68 8.67 4.61 4.21
Estonia 8.72 7.74 5.95 5.01 3.43 2.68
France 10.08 8.92 6.13 5.25 3.17 2.59
Germany 12.68 9.01 7.17 5.07 3.36 2.33
Ghana 10.47 9.01 6.89 6.01 4.01 3.17
Iceland 10.01 9.26 6.82 5.92 3.73 3.01
Iran, Islamic Republic of 8.18 6.92 6.01 4.72 3.57 2.46
Iraq 10.93 11.60 7.87 7.74 4.73 4.17
Ireland 15.65 11.93 9.01 6.67 4.01 2.91
Israel 12.01 9.18 7.36 5.47 3.81 2.67
Italy 9.18 7.51 5.17 4.22 2.45 2.00
Jamaica 20.03 20.10 12.60 13.35 6.19 6.59
Japan 16.58 10.25 10.28 6.01 4.52 2.50
Jordan 8.75 9.08 6.09 6.24 3.54 3.42
Kenya 10.06 10.60 7.25 7.01 4.15 4.00
Kuwait 12.04 11.35 8.81 7.50 4.89 4.01
Lebanon 11.93 11.01 7.09 6.84 3.61 3.17
Lithuania 6.67 7.09 4.92 4.98 3.00 2.77
Moldova, Republic of 10.09 10.43 7.01 7.01 4.01 4.01
Montenegro 10.52 10.01 7.01 6.84 4.01 3.59
Morocco 8.77 8.71 5.71 5.62 3.01 2.87
Netherlands 8.56 7.82 5.16 4.64 2.67 2.17
New Zealand 15.01 10.57 8.59 6.18 4.00 2.92
Nigeria 10.38 10.30 7.05 6.98 4.05 3.85
Norway 10.01 9.01 6.32 5.56 3.30 2.76
Oman 11.01 8.73 8.10 6.13 4.64 3.18
Philippines 14.36 13.10 9.87 8.53 5.62 4.42
Poland 9.01 8.39 5.67 5.11 3.01 2.83
Portugal 9.92 9.01 6.02 5.42 3.01 2.52
Romania 9.77 9.01 5.90 5.51 2.94 2.82
Russian Federation 10.01 11.30 6.47 6.92 3.34 3.25
Saudi Arabia 8.45 7.95 5.69 5.05 3.13 2.59
Serbia 9.51 9.84 5.84 6.01 3.00 3.01
South Africa 13.25 11.34 7.92 6.93 3.90 3.26
Spain 10.89 9.10 6.73 5.50 3.12 2.58
Sri Lanka 10.09 12.01 7.09 8.26 3.98 4.26
Sweden 10.01 7.76 6.05 4.64 3.06 2.18
Switzerland 11.31 8.79 6.68 5.01 3.21 2.34
Syrian Arab Republic 16.56 16.01 11.09 10.66 5.34 5.64
Tunisia 8.26 7.68 5.71 5.01 3.23 2.60
Türkiye 7.34 7.02 5.01 4.73 2.77 2.50
United Arab Emirates 10.75 8.01 6.92 5.13 3.67 2.58
United Kingdom 15.52 11.01 8.96 6.25 3.96 2.76
United States 14.27 11.84 8.51 7.01 4.01 3.09
Zimbabwe 12.01 10.76 8.11 7.44 4.84 4.22
Table S11: Spearman’s rank correlation between cumulative productivity and two conventional career metrics, total productivity (TP) and publishing-career length (PCL).
Female Male
Country TP PCL TP PCL
Albania 0.961 0.689 0.965 0.718
Algeria 0.975 0.636 0.973 0.698
Australia 0.957 0.677 0.937 0.662
Austria 0.951 0.672 0.924 0.657
Bahrain 0.986 0.504 0.969 0.533
Belarus 0.977 0.540 0.957 0.573
Belgium 0.908 0.648 0.882 0.619
Bosnia and Herzegovina 0.972 0.700 0.974 0.720
Bulgaria 0.976 0.677 0.971 0.663
Canada 0.958 0.644 0.934 0.639
Cyprus 0.970 0.700 0.947 0.730
Czechia 0.947 0.693 0.949 0.694
Denmark 0.957 0.699 0.933 0.677
Egypt 0.982 0.619 0.976 0.631
Estonia 0.957 0.723 0.944 0.687
France 0.949 0.688 0.932 0.670
Germany 0.957 0.643 0.924 0.639
Ghana 0.978 0.604 0.975 0.677
Iceland 0.957 0.698 0.952 0.719
Iran, Islamic Republic of 0.961 0.648 0.944 0.706
Iraq 0.981 0.539 0.979 0.571
Ireland 0.970 0.651 0.948 0.668
Israel 0.954 0.659 0.938 0.679
Italy 0.942 0.656 0.933 0.648
Jamaica 0.977 0.611 0.966 0.613
Japan 0.945 0.563 0.892 0.556
Jordan 0.973 0.566 0.963 0.673
Kenya 0.975 0.636 0.969 0.673
Kuwait 0.970 0.602 0.955 0.661
Lebanon 0.973 0.616 0.966 0.672
Lithuania 0.954 0.754 0.945 0.732
Moldova, Republic of 0.978 0.614 0.977 0.638
Montenegro 0.978 0.714 0.978 0.710
Morocco 0.969 0.616 0.971 0.678
Netherlands 0.945 0.693 0.921 0.669
New Zealand 0.967 0.674 0.948 0.680
Nigeria 0.975 0.552 0.973 0.608
Norway 0.960 0.715 0.944 0.694
Oman 0.973 0.485 0.949 0.652
Philippines 0.982 0.473 0.979 0.504
Poland 0.966 0.723 0.957 0.703
Portugal 0.965 0.705 0.951 0.687
Romania 0.965 0.710 0.960 0.716
Russian Federation 0.981 0.632 0.976 0.618
Saudi Arabia 0.979 0.567 0.966 0.662
Serbia 0.974 0.738 0.973 0.729
South Africa 0.971 0.676 0.960 0.675
Spain 0.954 0.666 0.939 0.661
Sri Lanka 0.971 0.598 0.977 0.607
Sweden 0.943 0.673 0.921 0.638
Switzerland 0.958 0.671 0.925 0.657
Syrian Arab Republic 0.986 0.439 0.978 0.484
Tunisia 0.954 0.666 0.945 0.698
Türkiye 0.961 0.721 0.949 0.725
United Arab Emirates 0.980 0.546 0.967 0.691
United Kingdom 0.964 0.646 0.938 0.639
United States 0.959 0.624 0.937 0.614
Zimbabwe 0.970 0.624 0.961 0.654
Table S12: Proportion of active authors in 2023 with a publication gap experience.
Country Female (%) Male (%)
Albania 7.87 9.97
Algeria 5.52 10.69
Australia 17.58 25.50
Austria 13.60 23.52
Bahrain 3.14 6.88
Belarus 5.03 9.77
Belgium 22.79 32.30
Bosnia and Herzegovina 7.97 12.04
Bulgaria 10.76 11.76
Canada 12.75 20.47
Cyprus 11.96 20.70
Czechia 17.25 23.76
Denmark 15.32 25.46
Egypt 6.09 8.51
Estonia 14.11 19.89
France 16.07 21.93
Germany 11.66 22.43
Ghana 4.41 8.64
Iceland 17.09 21.55
Iran, Islamic Republic of 9.08 19.17
Iraq 3.90 5.51
Ireland 12.83 19.97
Israel 11.66 19.21
Italy 22.12 28.37
Jamaica 6.31 7.62
Japan 13.04 29.59
Jordan 5.29 10.83
Kenya 6.83 9.34
Kuwait 6.35 11.52
Lebanon 7.12 12.55
Lithuania 16.29 19.46
Moldova, Republic of 6.63 8.67
Montenegro 9.26 14.43
Morocco 8.63 14.20
Netherlands 17.58 25.78
New Zealand 12.36 20.90
Nigeria 4.81 6.36
Norway 14.67 21.29
Oman 3.89 12.22
Philippines 2.85 4.08
Poland 15.56 21.78
Portugal 20.87 25.46
Romania 12.24 15.57
Russian Federation 7.34 8.40
Saudi Arabia 5.34 11.49
Serbia 16.08 17.08
South Africa 9.72 13.53
Spain 17.11 24.76
Sri Lanka 6.59 5.60
Sweden 20.20 29.48
Switzerland 13.23 22.33
Syrian Arab Republic 2.88 4.98
Tunisia 12.61 21.59
Türkiye 13.89 21.08
United Arab Emirates 5.10 13.53
United Kingdom 12.43 21.75
United States 12.48 19.49
Zimbabwe 5.55 8.66
Table S13: Proportion of returning authors within newly active authors in 2023.
Country Female (%) Male (%)
Albania 17.56 18.82
Algeria 7.35 11.07
Australia 16.87 24.22
Austria 12.51 25.09
Bahrain 1.66 6.82
Belarus 5.94 10.71
Belgium 24.22 34.90
Bosnia and Herzegovina 8.45 13.62
Bulgaria 9.76 10.83
Canada 12.86 22.26
Cyprus 13.08 20.73
Czechia 19.34 23.93
Denmark 14.37 25.84
Egypt 4.42 7.18
Estonia 15.22 21.03
France 17.21 22.53
Germany 11.27 23.46
Ghana 4.29 8.65
Iceland 18.56 19.40
Iran, Islamic Republic of 11.85 21.75
Iraq 4.20 6.28
Ireland 11.94 19.56
Israel 11.75 18.59
Italy 20.46 25.42
Jamaica 13.64 4.17
Japan 13.79 34.83
Jordan 3.66 9.64
Kenya 6.23 9.35
Kuwait 7.65 14.32
Lebanon 5.03 7.28
Lithuania 14.71 21.62
Moldova, Republic of 6.99 9.66
Montenegro 8.97 9.76
Morocco 5.56 9.28
Netherlands 16.32 25.55
New Zealand 13.87 20.33
Nigeria 5.22 8.06
[…]